OpenStudy (anonymous):

What is the simplified form of (x^2 + 5x +4)/(x - 4) multiplied by 3/(x - 1) divided by (x + 4)/(2x^2 + 2x - 4)?

OpenStudy (anonymous):

\[(x^2+5x+4)/(x-4)=(x+4)(x+1)/(x-4)\] \[3/(x-1)*(2x^2+2x-4)/(x+4) = 6(x+2)(x-1)/(x-1)(x+4)=6(x+2)/(x+4)\] \[(x+4)(x+1)/(x-4)*6(x+2)/(x+4) = 6(x+2)(x+1)/(x-4)\] \[6(x+2)(x+1)/(x-4)\]

OpenStudy (anonymous):

Thanks for your help!